# Scaling of diffusion constants in perturbed easy-axis Heisenberg spin chains

### v1

@@ Abstract
$\langle S^{z}_{r}(t)S_{r^{\prime}}^{z}(0)\rangle=\frac{\text{tr}[e^{iHt}S^{z}_{r}e^{-iHt}S_{r^{\prime}}^{z}]}{2^{N}}\;.$ (S17)
@@
$C_{r}(t)=\langle S^{z}_{r}(t)S_{1}^{z}(0)\rangle-\langle S^{z}_{r}(t)S_{N/2+1}^{z}(0)\rangle\,,$ (S18)
@@
$\langle S_{r}^{z}(t)\rangle\approx\frac{1}{T_{\text{max}}}\sum_{T=1}^{T_{\text{max}}}d_{r,T}(t)\,,$ (S20)
@@
$\tau_{j+1}=\tau_{j}-\frac{\ln(\varepsilon_{j+1})}{2\gamma}\,,$ (S21)
where $\varepsilon_{j+1}$ are random numbers drawn from a uniform distribution ]0,1]]0,1]] 0 , 1 ]. The amplitudes $A_{j}$ in Eq. (S19) read
$A_{j}=\frac{a_{j}\ -\ d_{1,T}(\tau_{j}-0^{+})}{\mu}$ (S22)
@@
Current in the steady state
Figure S7: Open boundary conditions and large system-bath coupling. TEBD data is compared to the results in Fig. 3 (b) of the main text.
@@
$\langle\delta S_{1}^{z}(t)\rangle\approx\frac{1}{T_{\text{max}}}\sum_{T=1}^{T_{\text{max}}}\delta d_{1,T}(t)$ (S24)
@@
$\delta d_{1,T}(t)=2\mu\sum_{j}A_{j}\Theta(t-\tau_{j})\ \langle[S_{1}^{z}(0)]^{2}\rangle\;.$ (S25)
@@
$\langle j_{r}\rangle=\langle j_{r^{\prime}}\rangle\,,\quad 1\leq r,r^{\prime}\leq N/2+1\,.$ (S26)
@@
$\langle j_{1}\rangle=\frac{\text{d}}{\text{d}t}\frac{\langle\delta S_{1}^{z}(t)\rangle}{2}\;,$ (S27)
@@
$D=-\frac{\langle j_{1}\rangle}{\langle S_{r+1}^{z}\rangle-\langle S_{r}^{z}\rangle}\;,$ (S28)
@@
Open boundary conditions
While we have focused throughout the main text on the case of periodic boundary conditions, it is certainly an interesting and important question whether and to what degree the results change for the case of open boundary conditions. Unfortunately, our approach to open systems is less reliable in this case. Instead, we show in Fig. S7 results from time-evolving block decimation (TEBD) for a large system-bath coupling $\gamma/J=1$ and compare to the results in Fig. 3 (b) of the main text.
@@
[…]

### v2

@@ B.1 Time evolution of densities
$\langle S^{z}_{r}(t)S_{r^{\prime}}^{z}(0)\rangle=\frac{\text{tr}[e^{iHt}S^{z}_{r}e^{-iHt}S_{r^{\prime}}^{z}]}{2^{N}}\;.$ (29)
@@
$C_{r}(t)=\langle S^{z}_{r}(t)S_{1}^{z}(0)\rangle-\langle S^{z}_{r}(t)S_{N/2+1}^{z}(0)\rangle\,$, (30)
@@
$\langle S_{r}^{z}(t)\rangle\approx\frac{1}{T_{\text{max}}}\sum_{T=1}^{T_{\text{max}}}d_{r,T}(t)\,$, (32)
@@
$\tau_{j+1}=\tau_{j}-\frac{\ln(\varepsilon_{j+1})}{2\gamma}\,$, (33)
where $\varepsilon_{j+1}$ are random numbers drawn from a uniform distribution ]0,1]]0,1]. The amplitudes $A_{j}$ in Eq. (31) read
$A_{j}=\frac{a_{j}\ -\ d_{1,T}(\tau_{j}-0^{+})}{\mu}$ (34)
@@
B.2 Current in the steady state
Figure 11: Open boundary conditions and large system-bath coupling. TEBD data is compared to the results in Fig. 3 (b) of the main text.
@@ B.2 Current in the steady state
$\langle\delta S_{1}^{z}(t)\rangle\approx\frac{1}{T_{\text{max}}}\sum_{T=1}^{T_{\text{max}}}\delta d_{1,T}(t)$ (36)
@@
$\delta d_{1,T}(t)=2\mu\sum_{j}A_{j}\Theta(t-\tau_{j})\ \langle[S_{1}^{z}(0)]^{2}\rangle\;.$ (37)
@@
$\langle j_{r}\rangle=\langle j_{r^{\prime}}\rangle\,,\quad 1\leq r,r^{\prime}\leq N/2+1\,.$ (38)
@@
$\langle j_{1}\rangle=\frac{\text{d}}{\text{d}t}\frac{\langle\delta S_{1}^{z}(t)\rangle}{2}\;,$ (39)
@@
$D=-\frac{\langle j_{1}\rangle}{\langle S_{r+1}^{z}\rangle-\langle S_{r}^{z}\rangle}\;,$ (40)
@@
B.3 Open boundary conditions
While we have focused throughout the main text on the case of periodic boundary conditions, it is certainly an interesting and important question whether and to what degree the results change for the case of open boundary conditions. Unfortunately, our approach to open systems is less reliable in this case. Instead, we show in Fig. 11 results from time-evolving block decimation (TEBD) for a large system-bath coupling $\gamma/J=1$ and compare to the results in Fig. 3 (b) of the main text.
@@ B.3 Open boundary conditions
Figure 12: The same data as in Fig. 4, but now depicted in a log-log plot and compared to the power law $\propto(1/N)^{-1/2}$.
Appendix C Details on the simple model
C.1 Time-dependent diffusion coefficient
In the main text, we have briefly introduced a simple model for the system-size scaling of the diffusion constant in the open system. Thus, we provide here an extended discussion and further details.
The simple model is based on the closed system and makes an assumption on the time-dependent diffusion coefficient in Eq. (6),
$D(t)=\frac{1}{\chi}\int_{0}^{t}\text{d}t^{\prime}\,\langle j(t^{\prime})j\rangle\,.$ (41)
Specifically, this assumption reads
$D(t)=at\,\Theta(t_{1}-t)+b\,\Theta(t-t_{1})+c\,(t-t_{2})\,\Theta(t-t_{2})\,$, (42)
where $\Theta(t)$ is the Heavyside function. Hence, $D(t)$ first increases linearly up to a time $t_{1}$, then remains constant up to a time $t_{2}$, and finally increases linearly again. In particular, as illustrated in Fig. 4, the quantities $t_{2}$ and $c$ are assumed to depend on system size as [58]
$t_{2}\propto N\,,\quad c\propto\frac{1}{N}\,$, (43)
while the other quantities $t_{1}$, $a$, and $b$ are independent of system size. The parameter $a$ is given by
$a=\frac{\langle j^{2}\rangle}{\chi}=\frac{1}{2}\,$ (44)
and the remaining parameters are obtained from a fit to $N=20$ data for $\Delta=1.5$,
$t_{1}=1.2\,,\quad t_{2}=0.5\,L\,,\quad b=0.63\,,\quad c=\frac{1.4}{N}\,.$ (45)
Note that the parameter $c$ results from a saturation of the current autocorrelation at long times,
$\lim_{t\to\infty}\frac{\langle j(t)j\rangle}{\chi}\propto\frac{1}{N}\,$, (46)
and plays the role of the finite-size Drude weight, which is expected to vanish for $\Delta=1.5$ and $N\to\infty$ [6]. Thus, for $N\to\infty$, the assumption in Eq. (42) leads to
$\lim_{t\to\infty}D(t)=b=\text{const.}\,$, (47)
which indicates normal diffusion, as expected for $\Delta=1.5$ [6]. It is worth pointing out that the finite-size Drude weight could in principle hide anomalous behavior,
$\frac{\langle j(t)j\rangle}{\chi}\propto\frac{1}{t}\,$, (48)
and therefore a logarithmic divergence of the diffusion coefficient,
$D(t)\propto\log t\,.$ (49)
However, numerical simulations like ours in Fig. 6 do not support such an anomalous contribution and already indicate a constant plateau. Moreover, for the following discussion, a potentially anomalous contribution plays a minor role. In fact, the finite-size Drude weight is much more important.
C.2 Real-time broadening
Next, let us stay in the closed system and consider the spatio-temporal correlation functions $\langle S^{z}_{r}(t)S_{r^{\prime}}^{z}(0)\rangle$ in Eq. (29). At $t=0$,
$\langle S^{z}_{r}(0)S_{r^{\prime}}^{z}(0)\rangle=\frac{1}{4}\,\delta_{r,r^{\prime}}$ (50)
at infinite temperature, where $\delta_{r,r^{\prime}}$ denotes the Kronecker symbol. Therefore, the profile is initially a sharp peak at the lattice site $r=r^{\prime}$. The real-time broadening of this peak can be followed via the spatial variance,
$\sigma^{2}(t)\!$ $=$ $\!\sum_{r}(r-r^{\prime})^{2}\,\frac{\langle S^{z}_{r}(t)S_{r^{\prime}}^{z}(0)\rangle}{1/4}$ (51)
$+$ $\!\Big{[}\sum_{r}(r-r^{\prime})\,\frac{\langle S^{z}_{r}(t)S_{r^{\prime}}^{z}(0)\rangle}{1/4}\Big{]}^{2}\,$,
which can be written in the form of Eq. (11),
$\sigma^{2}(t)=2\int_{0}^{t}\text{d}t^{\prime}\,D(t^{\prime})\,$, (52)
and is expressed in terms of the time-dependent diffusion coefficient $D(t)$ in Eq. (6). Note that this expression does not require the assumption for $D(t)$ in Eq. (42). Now, let us consider the time $t^{*}$, when the intially sharp peak has broadened over half of the system,
$3\sigma(t^{*})=\frac{N}{2}\,$, (53)
where we take three times the standard deviation for sake of simplicity. Using the assumption in Eq. (42), $t^{*}$ can then be calculated for any system size $N$. For large $N$, one finds the asymptotic behavior
$t^{*}\propto N^{3/2}$ (54)
and, in particular, $t^{*}\gg t_{2}$. For normal diffusion, as a side remark, one would have $t^{*}\propto N^{2}$.
C.3 Link to the open system
Eventually, let us link to the actual open system. To this end, we assume that, for the open system, (i) $t^{*}$ is the relevant time scale and that (ii) $D(t=t^{*})$, evaluated at $t=t^{*}$, is the corresponding diffusion coefficient. Note that these two assumptions are reasonable, in view of the correspondence between closed and open systems in and around Eq. (32).
In Fig. 4, we have shown the finite-size scaling of the so obtained $D(t^{*})$. On the one hand, $D(t^{*})$ agrees nicely with the $1/N$ scaling, as suggested by numerics for small system sizes, which serves as a consistency check. On the other hand, $D(t^{*})$ departs from a $1/N$ scaling for larger system sizes. To conclude on the asymptotic behavior, we show in Fig. 12 the same data as in Fig. 4, but now depicted in a log-log plot. Clearly,
$D(t^{*})\propto\sqrt{N}\,$, (55)
which can be also seen easily by plugging the asymptotic behavior of the time $t^{*}\propto N^{3/2}$ into the assumption for the diffusion coefficient in Eq. (42).
We should stress that the assumption in Eq. (42) is only applicable to the integrable point. For the case of a weak but nonvanishing perturbation, this assumption is only expected to hold up to a certain system size. Above this system size, the scaling of the Drude weight changes from a power law to an exponential decrease in system size, which is consistent with the eigenstate thermalization hypothesis [61] and in turn leads to a convergence of the diffusion coefficient w.r.t. system size.